OpenStudy (anonymous):

Find dy/dx for 5x^2-2xy+7y^2=0. I got dy/dx=(5x-y)/(x-7y) but the answers doesn't seem to match. A) (5x+7y)/2 B) (y-5x)/(7y-x) C) 10x+12y D) (5x-y)/(7y-x) E) None of these Which one is the answer?

zepdrix (zepdrix):

Differentiating Implicitly will give us,\[\large 10x-2y-2xy'+14yy'=0\]Moving the non y' junk to the other side gives us,\[\large 14yy'-2xy'=2y-10x\]Dividing a 2 out of everything,\[\large 7yy'-xy'=y-5x\]Factoring out a y',\[\large y'(7y-x)=y-5x\]Solving for y',\[\huge y'=\frac{y-5x}{7y-x}\] Check my steps, maybe you can find your mistake somewhere :D

zepdrix (zepdrix):

Oh oh oh oh i see the issue..

zepdrix (zepdrix):

Your top is off by a negative, and your bottom is also off by a negative. So your answer is equivalent to mine. Just factor a negative out of the top 2 terms, and a negative out of the bottom 2 terms.